OpenStudy (anonymous):

What is the factored form of the expression? s^2 - 81

OpenStudy (anonymous):

\[s ^{2} - 81\]

OpenStudy (luigi0210):

Difference of two perfect squares: (x-a)(x+a)=x^2-a^2

OpenStudy (unklerhaukus):

\[(a-b)(a+b)=a^2+ab-ba-b^2=(a^2-b^2)\]

OpenStudy (anonymous):

(s-9)(s+9)

OpenStudy (anonymous):

So what do use for x and a @Luigi0210 the answer choices are: a. (s-9)(s-9) b. (s-9)(s+11) c.(s-9)(s+9) d.(s+9) (s+9) I am learning toward D and I'm pretty sure it's not b

OpenStudy (luigi0210):

x represents your variable a is your number

OpenStudy (luigi0210):

what two numbers multiply to 81?

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

9 and 9 that's why I was thinking definitely not b

OpenStudy (anonymous):

factor (s-9)(s+9) s^2 +9s-9s-81 (factor) s^2 -81 (simplify)

OpenStudy (anonymous):

Thanks

OpenStudy (luigi0210):

Yup :P Now, they need to have opposite signs so they can cancel out like drotz showed

OpenStudy (anonymous):

yeah so you guys should check out my question... its a brain stomper

OpenStudy (luigi0210):

link?

OpenStudy (anonymous):

http://openstudy.com/study#/updates/51aecad5e4b06ee3ee265fd2 i think that will work i hope